OpenStudy (babynini):

Find length and direction of u u = <-4 (sqrt3),4>

OpenStudy (babynini):

The initial point is (0,0)

OpenStudy (babynini):

@jim_thompson5910 :)

jimthompson5910 (jim_thompson5910):

u = <a,b> length of u = |u| = sqrt(a^2+b^2) this is the same as saying "the magnitude of vector u"

jimthompson5910 (jim_thompson5910):

direction: theta = arctan(b/a) if theta is negative, and you know for a fact that the vector is in Q2, then add on 180 degrees

OpenStudy (babynini):

oh, I didn't make that connection! ok.

OpenStudy (babynini):

and I didn't know that about direction at all xD thanks :) let me get the answers real quick.

OpenStudy (babynini):

Length = 8 ?

OpenStudy (babynini):

Direction = 178.95

jimthompson5910 (jim_thompson5910):

I agree with the length, but not the direction

jimthompson5910 (jim_thompson5910):

b/a = 4/(-4*sqrt(3)) = -0.57735026918962 arctan(b/a) = arctan(-0.57735026918962) = -29.9999999999998 which is essentially -30 add on 180 -30+180 = 150

OpenStudy (babynini):

hrm when I do = arctan (4/(-4 sqrt3)) it comes out to = -0.52

jimthompson5910 (jim_thompson5910):

you must be in radian mode

OpenStudy (babynini):

...oh xD

OpenStudy (babynini):

Yeah you're right, I was.

jimthompson5910 (jim_thompson5910):

if you're using desmos, then click the wrench to change to degree mode

OpenStudy (babynini):

Scientific calculator :) thanks though, I use desmos as well.